For the function g(x) = x2 + 5x - 2, determine g(-3x). Thank you
Answer by Theo(13342)

the function is g(x) = x^2 + 5x - 2

to find g(-3x), then replace x with -3x in the g(x) function to get:

g(-3x) = (-3x)^2 + 5*(-3x) - 2

simplify to get g(-3x) = 9x^2 -15x - 2

to determine whether you did this right, let x equal a random value of your choosing.

i chose 15.

g(-3x) would then be equal to g(-45).

g(x) = x^2 + 5x - 2 becomes g(-45) = (-45)^2 + 5*-45 - 2 which becomes 2025 - 225 - 2 which is equal to 1798.

the function of g(-3x) can be graphed as shown below.

g(x) = x^2 + 5x - 2
g(-3x) = (-3x)^2 + 5*(-3x) - 2

here's what the graph looks like.

in the graph, y = g(x).